OpenStudy (anonymous):

how do i factor (5a-2b)(5a-2b)

OpenStudy (blurbendy):

5a * 5a = 25a^2 5a * -2b = -10ab -2b * 5a = -10ab -2b * - 2b = 4b^2 So, you end up with 25a^2 - 20ab + 4b^2 Can you finish it from here?